# Understanding Deep Representation Learning via Layerwise Feature Compression and Discrimination

### v3

@@ 1.3 Notation & Paper Organization
[…]
@@
The rest of the paper is organized as follows. In Section 2, we introduce the basic settings of our problem. In Section 3, we describe the main result of our work and discuss its implications, with proofs provided in Section B in the appendix. In Section 4, we discuss the relationship of our results to related works. In Section 5, we verify our theoretical claims and investigate nonlinear networks via numerical experiments. Finally, we draw a conclusion and discuss future directions in Section 6. We defer all the auxiliary technical results to the appendix.
@@ 2 Preliminaries
In this section, we first formally set up the problem of training DLNs for solving multi-class classification problems in Section 2.1, followed by introducing the metrics for measuring within-class compression and between-class discrimination of features at each layer in Section 2.2.
@@ 2.1 Problem Setup
We consider a $K$-class classification problem with training samples and labels $\{(\bm{x}_{k,i},\bm{y}_{k})\}_{i\in[n_{k}],k\in[K]}$, where $\bm{x}_{k,i}\in\mathbb{R}^{d}$ is the $i$-th sample in the $k$-th class, and $\bm{y}_{k}\in\mathbb{R}^{K}$ is an one-hot label vector with the $k$-th entry being $1$ and 00 elsewhere. We denote by $n_{k}$ the number of samples in the $k$-th class for each $k\in[K]$. Here, we assume that the number of samples in each class is the same, i.e., $n_{1}=\cdots=n_{K}=n$. Moreover, we denote the total number of samples by $N=nK$. Without loss of generality, we can arrange the training samples in a class-by-class manner such that
@@
where $\otimes$ denotes the Kronecker product. Note that we also use $\bm{x}_{i}$ to denote the $i$-th column of $\bm{X}$.
@@
where $\bm{W}_{1}\in\mathbb{R}^{d_{1}\times d}$, $\bm{W}_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}\;(l=2,\dots,L-1)$, and $\bm{W}_{L}\in\mathbb{R}^{K\times d_{L-1}}$ are weight matrices. It is worth mentioning that the last-layer weight $\bm{W}_{L}$ is referred to as the linear classifier and the output of the $l$-th layer as the $l$-th layer feature for all $l\in[L-1]$ in the literature. As discussed in Section 1, DLNs are often used as prototypes for studying practical deep networks [50, 65, 73, 77, 99]. Then, we train an $L$-layer linear network to learn weights $\bm{\Theta}=\{\bm{W}_{l}\}_{l=1}^{L}$ via minimizing the mean squared error (MSE) loss between $f_{\bm{\Theta}}(\bm{x})$ and $\bm{y}$, i.e.,
@@
  • First, the cross-entropy (CE) loss is arguably the most popular loss function used to train neural networks for classification problems [127]. However, recent studies [47, 128] have demonstrated through extensive experiments that the MSE loss achieves comparable or even superior performance than the CE loss across various tasks.
  • Second, while DLNs may appear simple, their loss landscapes, as defined in (3), are highly nonconvex, leading to highly nonlinear learning dynamics in gradient descent (GD). Notably, the nonlinear GD dynamics in DLNs closely mirror that in their nonlinear counterparts [75, 99, 100].
  • Finally, DLNs tend to be over-parameterized, with width of networks $d_{l}$ and feature dimensions $d$ exceeding the training samples $N$. In this context, Problem (3) has infinitely many solutions that can achieve zero training loss, i.e., $\bm{W}_{L:1}\bm{X}=\bm{Y}$. However, many studies have delved into the convergence behavior of GD by closely examining its learning trajectory. These investigations reveal that GD, when initialized appropriately, exhibits an implicit bias towards minimum norm solutions [14, 86] with approximately balanced and low-rank weights [2, 86], which will be discussed further Section 3.2.
@@ 2.2 The Metrics of Feature Compression and Discrimination
In this work, we focus on studying feature structures at each layer in a trained DLN. Given the weights $\bm{\Theta}=\{\bm{W}_{l}\}_{l=1}^{L}$ satisfying $\bm{W}_{L:1}\bm{X}=\bm{Y}$, by construction the weights $\{\bm{W}_{l}\}_{l=1}^{L}$ of the DLN transform the input data $\bm{X}$ into the membership matrix $\bm{Y}$ at the final layer. However, the hierarchical structure of the DLN prevents us from gaining insight into the underlying mechanism of how it transforms the input data into output from shallow to deep layers. To unravel this puzzle, we probe the features learned at intermediate layers. In our setting, we write the $l$-th layer’s feature of an input sample $\bm{x}_{k,i}$ as
@@
$\bm{\mu}_{k}^{l}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\bm{z}_{k,i}^{l},\quad\bm{\mu}^{l}=\frac{1}{K}\sum_{k=1}^{K}\bm{\mu}_{k}^{l}$ (7)
@@
Definition 1.
@@
$\boxed{\quad C_{l}=\frac{\mathrm{Tr}(\bm{\Sigma}_{W}^{l})}{\mathrm{Tr}(\bm{\Sigma}_{B}^{l})},\quad D_{l}=1-\max_{k\neq k^{\prime}}\frac{\langle\bm{\mu}_{k}^{l},\bm{\mu}_{k^{\prime}}^{l}\rangle}{\|\bm{\mu}_{k}^{l}\|\|\bm{\mu}_{k^{\prime}}^{l}\|}\quad}$ (8)
are the metrics of the within-class compression and between-class discrimination of intermediate features at the $l$-th layer, respectively.
@@
The study of feature compression has recently caught great attention in both supervised [123, 38, 91] and unsupervised [109, 106] deep learning. For our definition of feature compression in (8), intuitively the numerator $\mathrm{Tr}(\bm{\Sigma}_{W}^{l})$ of the metric $C_{l}$ measures how well the features from the same class are compressed towards the class mean at the $l$-th layer. More precisely, the features of each class are more compressed around their respective means as $\mathrm{Tr}(\bm{\Sigma}_{W}^{l})$ decreases. On the other hand, the denominator $\mathrm{Tr}(\bm{\Sigma}_{B}^{l})$ serves as a normalization factor, rendering this metric invariant to the scale of features. Specifically, given some weights $\bm{W}_{1},\dots,\bm{W}_{l}$ with $C_{l}=\mathrm{Tr}(\bm{\Sigma}_{W}^{l})/\mathrm{Tr}(\bm{\Sigma}_{B}^{l})$, if we scale them to $t\bm{W}_{1},\dots,t\bm{W}_{l}$ for some $t>0$, then $C_{l}$ does not change, while the corresponding numerator becomes $t^{l}\mathrm{Tr}(\bm{\Sigma}_{W}^{l})$. It should be noted that this metric and similar metrics have been studied in prior works [70, 96, 112, 124]. For instance, [112] employed this metric to measure the variability of within-class features to simplify theoretical analysis. Moreover, a similar metric $\mathrm{Tr}(\bm{\Sigma}_{W}^{l}\bm{\Sigma}_{B}^{l^{\dagger}})$ has been used to characterize within-class variability collapse in recent studies on neural collapse in terms of last-layer features [38, 91, 96, 125, 127]. Our studied metric $C_{l}$ can be viewed as its simplification. Additionally, [53] employed $\mathrm{Tr}(\bm{\Sigma}_{W}^{l}\bm{\Sigma}_{B}^{l^{\dagger}})$ to measure how well the data are separated across intermediate layers. Due to their similarity, $C_{l}$ can also serve as a metric for measuring data separation.
@@
It is worth pointing out that learning discriminative features have a long history back to unsupervised dictionary learning [29, 6, 103, 94, 129]. For our definition of feature discrimination, $\arccos\left(\langle\bm{\mu}_{k}^{l},\bm{\mu}_{k^{\prime}}^{l}\rangle/(\|\bm{\mu}_{k}^{l}\|\|\bm{\mu}_{k^{\prime}}^{l}\|)\right)$ computes the angle between class means $\bm{\mu}_{k}^{l}$ and $\bm{\mu}_{k^{\prime}}^{l}$. This, together with (8), indicates that $D_{l}$ measures the feature discrimination by calculating the smallest angles among feature means of all pairs. Moreover, we can equivalently rewrite $D_{l}$ in (8) as
@@
This indicates that $D_{l}$ computes the smallest distance between normalized feature means. According to these two interpretations, features between classes become more discriminative as $D_{l}$ increases. Recently, [83] considered a variant of the inter-class variance $\sum_{k=1}^{K}\sum_{k^{\prime}\neq k}\|\bm{\mu}_{k}-\bm{\mu}_{k^{\prime}}\|^{2}$ to measure linear separability of representations of deep networks.
@@ 3 Main Results
In this section, we present our main theoretical result based on the problem setup and metrics introduced in Section 2. Specifically, we describe our main theory in Section 3.1, followed by a discussion of assumptions in Section 3.2.
@@ 3.1 Main Theorem
Before we present the main theorem, we make the following assumptions on the data input and the weights of the DLN in (2).
Assumption 1.
@@
$\left|\|\bm{x}_{i}\|^{2}-1\right|\leq\frac{\theta}{N},\ \left|\langle\bm{x}_{i},\bm{x}_{j}\rangle\right|\leq\frac{\theta}{N},\ \forall 1\leq i\neq j\leq N,$ (9)
@@
Discussion on 1.
Remark that we make this assumption to simplify our analysis. It can be relaxed by enhancing the analysis, and it might even be violated in empirical scenarios. Here, the condition $d\geq N$ guarantees that $\bm{X}$ is linearly separable in the sense that there exists a linear classifier $\bm{W}\in\mathbb{R}^{K\times d}$ such that $\bm{W}\bm{x}_{k,i}=\bm{y}_{k}$ for all $i,k$. It is worth pointing out that the same condition has been studied in [22, 23, 39], and similar linear separability conditions have been widely used for studying implicit bias of gradient descent [89, 92, 98]. Notably, this condition remains valid in nonlinear networks in the sense that the intermediate features generated by the initial layers exhibit linear separability as shown in [1, 7, 83, 95] (see Figure 1, where the near-perfect accuracy at intermediate layers shows that the features at that layer are already linearly separable.). In addition, nearly orthonormal data is commonly used in the theoretical analysis of learning dynamics for training neural networks; see, e.g., [15, 39, 92]. In particular, this condition holds with high probability when dealing with well-conditioned Gaussian distributions, as indicated in [39]. It generally applies to a broad class of subgaussian distributions, as demonstrated in Claim 3.1 of [54].
@@
Since the DLN in Problem (3) is over-parameterized, it has infinitely many solutions satisfying $\bm{W}_{L:1}\bm{X}=\bm{Y}$. However, GD for training networks typically has an implicit bias towards certain solutions with benign properties [4, 46, 64, 86, 105, 98]. Specifically, previous work has demonstrated that, with assumptions concerning network initialization and the dataset, gradient flow tends to favor solutions with minimum norms and balanced weights; see, e.g., [86, 22, 3, 31]. Recent studies have also revealed that GD primarily updates a minimal invariant subspace of the weight matrices, thereby preserving the approximate low-rankness of the weights across all layers [51, 124]. Based on these observations, we assume that the trained weights $\bm{\Theta}$ satisfy the following benign properties to investigate how trained deep networks hierarchically transform input data into labels.
Assumption 2.
@@
We defer the discussion on 2 to Section 3.2, where we provide both theoretical and empirical evidence to support it. It is worth noting that 2 is intended as an initial framework for understanding progressive feature behavior and can be relaxed into more general conditions. For example, a follow-up work [62, Theorem 3.1] proves related but narrower results with weakened versions of the conditions in 2 (e.g. (11) holds only approximately). Building upon the above assumptions, we now present our main theorem concerning hierarchical representations in terms of feature compression and discrimination.
Theorem 1.
@@
(i) Progressive within-class feature compression: It holds that
@@
(ii) Progressive between-class feature discrimination: For all $l\in[L-1]$, we have
@@
(a) $6$-layer linear network
(b) $10$-layer nonlinear network
Figure 4: Progressive feature compression and discrimination on both linear and nonlinear networks. We plot the feature compression and discrimination metrics defined in (8) at each layer for both the linear network (top row) and nonlinear network (bottom row). We train both networks using a nearly orthogonal dataset as described in 1, initializing the network weights satisfying (18), with an initialization scale of $\xi=0.3$. We train both networks via gradient descent until convergence. In each figure, the $x$-axis denotes the number of layers from shallow to deep. In the left figures, the $y$-axis denotes the compression measure $C_{l}$ in the logarithmic scale; In the right figures, the $y$-axis denotes the discrimination measure $D_{l}$. More experimental details can be found in Section 5.2.2.
We defer the detailed proof to Appendix B. As we discussed in Section 1, numerous empirical studies have been conducted to explore the feature structures of intermediate and final layers in deep networks, particularly concerning feature compression and discrimination; see, e.g., [1, 7, 17, 25, 35, 43, 53, 48, 83, 96, 95, 126]. However, there remains a scarcity of theoretical analysis to elucidate their observations. Despite the acknowledged limitations of DLNs, our work takes the first step towards rigorously developing a unified framework to quantify feature compression and discrimination across different layers. Specifically, 1 shows that, given input data satisfying 1, for a trained DLN with weights satisfying 2, features within the same class are compressed at a geometric rate on the order of ${\varepsilon^{2}}/{n^{1/L}}$ when $n$ is sufficiently large, while the features between classes are discriminated at a linear rate on the order of $1/L$ across layers. These findings are strongly supported by our empirical results in Figure 4 (top figures), and also appear on nonlinear networks, as shown in Figure 4 (bottom figures). In the following, we discuss the implications of our main result.
@@
Although our result is rooted in DLNs, it provides valuable insights into the feature evolution in nonlinear networks. Specifically, the linear separability of features learned by initial layers in nonlinear networks (see Figure 1) allows subsequent layers to be effectively replaced by linear counterparts for compressing within-class features and discriminating between-class features (see Figure 5). Therefore, studying DLNs helps us understand the role of the nonlinear layers after the initial layers in nonlinear networks for learning features. This understanding also sheds light on the pattern where within-class features compress at a geometric rate and between-class features discriminate at a linear rate in nonlinear networks, as illustrated in Figure 4 (bottom figures). A natural direction is to extend our current analysis framework to nonlinear networks, especially homogeneous neural networks [78].

### v4

@@ 1.2 Notation and Paper Organization
Let $\mathbb{R}^{n}$ be the $n$-dimensional Euclidean space and $\|\cdot\|$ be the Euclidean norm. Given a matrix $\bm{A}\in\mathbb{R}^{m\times n}$, we use $\bm{a}_{i}$ to denote its $i$-th column; we use $\|\bm{A}\|_{F}$ to denote the Frobenius norm of $\bm{A}$; we use $\sigma_{\max}(\bm{A})$ (or $\|\bm{A}\|$), $\sigma_{i}(\bm{A})$, and $\sigma_{\min}(\bm{A})$ to denote the largest, the $i$-th largest, and the smallest singular values, respectively. We use $\bm{A}^{\dagger}$ to denote the pseudo-inverse of a matrix $\bm{A}$. Given $L\in\mathbb{N}$, we use $[L]$ to denote the index set $\{1,\cdots,L\}$. Let $\mathcal{O}^{n}=\{\bm{Z}\in\mathbb{R}^{n\times n}:\bm{Z}^{T}\bm{Z}=\bm{I}\}$ denote the set of all $n\times n$ orthogonal matrices. We denote the Kronecker product by $\otimes$. Given weight matrices $\bm{W}_{1},\dots,\bm{W}_{L}$, let $\bm{W}_{l:1}:=\bm{W}_{l}\cdots\bm{W}_{1}$ denote a matrix multiplication from $\bm{W}_{l}$ to $\bm{W}_{1}$ for all $l\in[L]$.
@@
The rest of the paper is organized as follows. In Section 2, we introduce the basic problem setup. In Section 3, we present the main results and discuss their implications, with proofs provided in Section A in the appendix. In Section 4, we discuss the connections of our results to related works. In Section 5, we validate our theoretical claims and investigate nonlinear networks through numerical experiments. Finally, we conclude and discuss future directions in Section 6. We defer all the auxiliary technical results to the appendix.
@@ 2 Preliminaries
In this section, we first formally introduce the problem of training DLNs for solving multi-class classification problems in Section 2.1, and then present the metrics for measuring within-class compression and between-class discrimination of features at each layer in Section 2.2.
@@ 2.1 Problem Setup
We consider a $K$-class classification problem with training samples and labels $\{(\bm{x}_{k,i},\bm{y}_{k})\}_{i\in[n_{k}],k\in[K]}$, where $\bm{x}_{k,i}\in\mathbb{R}^{d}$ is the $i$-th sample in the $k$-th class, and $\bm{y}_{k}\in\mathbb{R}^{K}$ is an one-hot label vector with the $k$-th entry being 11 and 0 elsewhere. We denote by $n_{k}$ the number of samples in the $k$-th class for each $k\in[K]$. Here, we assume that the number of samples in each class is the same, i.e., $n_{1}=\cdots=n_{K}=n$. This assumption is commonly used in recent studies on deep representation learning for classification problems (zhu2021geometric; yaras2022neural; zhou2022optimization; zhou2022all). Moreover, we denote the total number of samples by $N=nK$. Without loss of generality, we arrange the training samples in a class-by-class manner such that
@@
where $\otimes$ denotes the Kronecker product. For convenience, we also use $\bm{x}_{i}$ to denote the $i$-th column of $\bm{X}$.
@@
where $\bm{W}_{1}\in\mathbb{R}^{d_{1}\times d}$, $\bm{W}_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ for $l=2,\dots,L-1$, and $\bm{W}_{L}\in\mathbb{R}^{K\times d_{L-1}}$ are the weight matrices. The last-layer weight $\bm{W}_{L}$ is referred to as the linear classifier and the output of the $l$-th layer as the $l$-th layer feature for all $l\in[L-1]$. As discussed in Section 1, DLNs are often used as prototypes for studying practical deep networks (hardt2016identity; kawaguchi2016deep; laurent2018deep; lu2017depth). We train an $L$-layer linear network to learn weights $\bm{\Theta}=\{\bm{W}_{l}\}_{l=1}^{L}$ via minimizing the mean squared error (MSE) loss between $f_{\bm{\Theta}}(\bm{x})$ and $\bm{y}$ over the training data, i.e.,
@@
  • First, the cross-entropy (CE) loss is arguably the most popular loss function used to train neural networks for classification problems (zhu2021geometric). However, recent studies (hui2020evaluation; zhou2022optimization) have demonstrated through extensive experiments that the MSE loss achieves performance comparable to or even superior to that of the CE loss across various tasks.
  • Second, while DLNs may appear simple, the loss landscape of the objective function as defined in (3) is highly nonconvex, leading to highly nonlinear learning dynamics in gradient descent (GD) (chen2025complete; achour2024loss). Notably, the nonlinear GD dynamics in DLNs closely mirror that in their nonlinear counterparts (lampinen2018analytic; saxe2019mathematical).
  • Finally, DLNs tend to be over-parameterized, with width of networks $d_{l}$ and feature dimensions $d$ exceeding the training samples $N$. In this setting, Problem (3) has infinitely many solutions that can achieve zero training loss, i.e., $\bm{W}_{L:1}\bm{X}=\bm{Y}$. Nevertheless, many studies have examined the convergence behavior of GD by closely examining its learning trajectory, revealing that GD—when initialized appropriately—exhibits an implicit bias towards minimum norm solutions (bartlett2020benign; min2021explicit) with approximately balanced and low-rank weights (arora2018convergence; min2021explicit), which will be discussed further in Section 3.2.
@@ 2.2 The Metrics of Feature Compression and Discrimination
In this work, we focus on studying feature structures at each layer in a trained DLN. Given the weights $\bm{\Theta}=\{\bm{W}_{l}\}_{l=1}^{L}$ satisfying $\bm{W}_{L:1}\bm{X}=\bm{Y}$, the weights $\{\bm{W}_{l}\}_{l=1}^{L}$ of the DLN transform the input data $\bm{X}$ into the membership matrix $\bm{Y}$ at the final layer. However, the hierarchical structure of the DLN prevents us from gaining insight into the underlying mechanism of how it transforms the input data into output from shallow to deep layers. To unravel this puzzle, we probe the features learned at intermediate layers. In our setting, we write the $l$-th layer’s feature of an input sample $\bm{x}_{k,i}$ as
@@
$\bm{\mu}_{k}^{l}:=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\bm{z}_{k,i}^{l},\quad\bm{\mu}^{l}:=\frac{1}{K}\sum_{k=1}^{K}\bm{\mu}_{k}^{l}$ (7)
@@
Definition 1 (Intermediate layer-wise feature compression and discrimination)
@@
$\quad C_{l}=\frac{\mathrm{Tr}(\bm{\Sigma}_{W}^{l})}{\mathrm{Tr}(\bm{\Sigma}_{B}^{l})}\quad\text{and}\quad D_{l}=1-\max_{k\neq k^{\prime}}\frac{\langle\bm{\mu}_{k}^{l},\bm{\mu}_{k^{\prime}}^{l}\rangle}{\|\bm{\mu}_{k}^{l}\|\|\bm{\mu}_{k^{\prime}}^{l}\|}$ (8)
are the metrics of within-class compression and between-class discrimination of intermediate features at the $l$-th layer, respectively.
@@
The study of feature compression has recently caught great attention in both supervised (yu2020learning; fang2021exploring; papyan2020prevalence) and unsupervised (shwartz2023compress) deep learning. For our definition of feature compression in (8), the numerator $\mathrm{Tr}(\bm{\Sigma}_{W}^{l})$ of the metric $C_{l}$ measures how well the features from the same class are compressed towards the class mean at the $l$-th layer. More precisely, the features of each class are more compressed around their respective means as $\mathrm{Tr}(\bm{\Sigma}_{W}^{l})$ decreases. The denominator $\mathrm{Tr}(\bm{\Sigma}_{B}^{l})$ prevents reporting spuriously small values of $C_{l}$ in near-collapse cases where all features approach a single mean, while simultaneously serving as a normalization factor that renders the metric invariant to feature scaling. Specifically, given some weights $\bm{W}_{1},\dots,\bm{W}_{l}$ with $C_{l}=\mathrm{Tr}(\bm{\Sigma}_{W}^{l})/\mathrm{Tr}(\bm{\Sigma}_{B}^{l})$, if we scale them to $t\bm{W}_{1},\dots,t\bm{W}_{l}$ for some $t>0$, then $C_{l}$ does not change, while the corresponding numerator becomes $t^{l}\mathrm{Tr}(\bm{\Sigma}_{W}^{l})$. It should be noted that this metric and similar ones have been studied in recent works (kothapalli2023neural; rangamani2023feature; tirer2023perturbation; yaras2023law). For instance, tirer2023perturbation employed this metric to measure the variability of within-class features to simplify theoretical analysis. Moreover, a similar metric $\mathrm{Tr}(\bm{\Sigma}_{W}^{l}\bm{\Sigma}_{B}^{l^{\dagger}})$ has been used to characterize within-class variability collapse in recent studies on neural collapse in terms of last-layer features (fang2021exploring; papyan2020prevalence; rangamani2023feature; yaras2022neural; zhu2021geometric). Our studied metric $C_{l}$ can be viewed as its simplification. Additionally, he2023law employed $\mathrm{Tr}(\bm{\Sigma}_{W}^{l}\bm{\Sigma}_{B}^{l^{\dagger}})$ to measure how well the data are separated across intermediate layers. Due to their similarity, $C_{l}$ can also serve as a metric for measuring data separation.
@@
It is worth pointing out that learning discriminative features has a long history, tracing back to unsupervised dictionary learning (donoho2003optimally; arora2014new; sun2016complete; Qu2020Geometric; zhai2020complete). For our definition of feature discrimination, $\arccos\left(\langle\bm{\mu}_{k}^{l},\bm{\mu}_{k^{\prime}}^{l}\rangle/(\|\bm{\mu}_{k}^{l}\|\|\bm{\mu}_{k^{\prime}}^{l}\|)\right)$ computes the angle between class means $\bm{\mu}_{k}^{l}$ and $\bm{\mu}_{k^{\prime}}^{l}$. This, together with (8), indicates that $D_{l}$ measures the feature discrimination by calculating the smallest angles among feature means of all pairs. Moreover, we can equivalently rewrite $D_{l}$ in (8) as
@@
This indicates that $D_{l}$ computes the smallest distance between normalized feature means. According to these two interpretations, features between classes become more discriminative as $D_{l}$ increases. Recently, masarczyk2023tunnel considered a variant of the inter-class variance $\sum_{k=1}^{K}\sum_{k^{\prime}\neq k}\|\bm{\mu}_{k}-\bm{\mu}_{k^{\prime}}\|^{2}$ to measure linear separability of representations of deep networks.
@@ 3 Main Results
In this section, we present our main theoretical results based on Section 2, first describing the theorem in Section 3.1 and then discussing the assumptions in Section 3.2.
@@ 3.1 Main Theorem
Before we present the main theorem, we make the following assumptions on the input data and the weights of the DLN in (2).
Assumption 2
@@
$\left|\|\bm{x}_{i}\|^{2}-1\right|\leq\frac{\theta}{N},\ \left|\langle\bm{x}_{i},\bm{x}_{j}\rangle\right|\leq\frac{\theta}{N},\ \text{for all}\ 1\leq i\neq j\leq N,$ (9)
@@
Discussion on Assumption 2.
We make this assumption primarily to simplify our analysis. It can be relaxed with a more refined analysis, and in practice, it may even be violated in empirical scenarios. Here, the condition $d\geq N$ guarantees that $\bm{X}$ is linearly separable in the sense that there exists a linear classifier $\bm{W}\in\mathbb{R}^{K\times d}$ such that $\bm{W}\bm{x}_{k,i}=\bm{y}_{k}$ for all $i,k$. We should point out that the same condition has been studied in (chatterji2023deep; chatterji2022interplay; frei2023implicit), and similar linear separability conditions have been widely used for studying implicit bias of gradient descent (nacson2019convergence; phuong2020inductive; soudry2018implicit). Notably, this condition also holds for nonlinear networks in the sense that the intermediate features generated by the initial layers exhibit linear separability as shown in alain2016understanding; ansuini2019intrinsic; masarczyk2023tunnel; recanatesi2019dimensionality (see Figure 1, where the near-perfect accuracy at intermediate layers shows that the features at that layer are already linearly separable.). In addition, nearly orthonormal data is commonly used in the theoretical analysis of learning dynamics for training neural networks; see, e.g., boursier2022gradient; frei2023implicit; phuong2020inductive. In particular, this condition holds with high probability for well-conditioned Gaussian distributions (frei2023implicit), and it generally applies to a broad class of subgaussian distributions, as demonstrated in Claim 3.1 of hu2020surprising.
@@
Since the DLN in Problem (3) is over-parameterized, it has infinitely many solutions satisfying $\bm{W}_{L:1}\bm{X}=\bm{Y}$. However, GD for training networks typically has an implicit bias towards certain solutions with benign properties (arora2019implicit; Gunasekar2017; Ji2019; min2021explicit; Shah2020; soudry2018implicit). In particular, prior work has demonstrated that, with assumptions on network initialization and the dataset, gradient flow tends to favor solutions with minimum norms and balanced weights; see, e.g., min2021explicit; chatterji2023deep; arora2018optimization; du2018algorithmic. Recent studies also reveal that GD primarily updates a minimal invariant subspace of the weight matrices, thereby preserving the approximate low-rankness of the weights across all layers (huh2021low; yaras2023law). Based on these findings, we assume that the trained weights $\bm{\Theta}$ satisfy the following benign properties to investigate how trained deep networks hierarchically transform input data into labels.
Assumption 3
@@
We defer the discussion of Assumption 3 to Section 3.2, where we provide both theoretical and empirical evidence in support of it. It is worth noting that Assumption 3 serves as an initial framework for understanding progressive feature behavior and can be relaxed to more general conditions. For example, a follow-up work (jacot2025wide, Theorem 3.1) proves related but narrower results under weakened versions of the conditions in Assumption 3 (e.g., (11) holds only approximately). Building upon the above assumptions, we now present our main theorem on hierarchical representations in terms of feature compression and discrimination.
Theorem 4
@@
(i) Progressive within-class feature compression: For $C_{l}$ in Definition 1, it holds that
@@
(ii) Progressive between-class feature discrimination: For $D_{l}$ in Definition 1, for all $l\in[L-1]$, we have
@@
We defer the detailed proof to Appendix A. As we discussed in Section 1, numerous empirical studies have been conducted to explore the feature structures of intermediate and final layers in deep networks, particularly concerning feature compression and discrimination; see, e.g., alain2016understanding; ansuini2019intrinsic; ben2022nearest; chen2022layer; galanti2022implicit; he2023law; hui2022limitations; masarczyk2023tunnel; rangamani2023feature; recanatesi2019dimensionality; zhang2022all. However, there remains a scarcity of theoretical analysis to elucidate their observations. Despite the acknowledged limitations of DLNs, our work takes the first step towards rigorously developing a unified framework to quantify feature compression and discrimination across different layers. Specifically, Theorem 4 shows that, given input data satisfying Assumption 2 and a trained DLN with weights satisfying Assumption 3, features within the same class are compressed at a geometric rate on the order of ${\varepsilon^{2}}/{n^{1/L}}$ when $n$ is sufficiently large, while the features between classes are discriminated at a linear rate on the order of $1/L$ across layers. In Figure 5, we also show empirical support for these findings in Figure 5 (top figures), where we simulate data and train a linear network according to our assumptions and plot the metrics at each layer. This phenomenon also appears in nonlinear networks, as shown in Figure 5 (bottom figures), where we train a nonlinear MLP with the same synthetic training data and plot the metrics at each layer. In the following, we discuss the implications of our main result.
(a) A 66-layer linear network
(b) A 1010-layer nonlinear network
Figure 5: Progressive feature compression and discrimination on both linear and nonlinear networks. We plot the feature compression and discrimination metrics defined in (8) for $l=1,\dots,L-1$ on both the linear network (top row) and nonlinear network (bottom row). We train both networks using a nearly orthogonal dataset as described in Assumption 2, initializing the network weights satisfying (18), with an initialization scale of $\xi=0.3$. We train both networks via gradient descent until convergence. In each figure, the $x$-axis denotes the number of layers from shallow to deep, with layer-0 denoting the inputs. In the left figures, the $y$-axis denotes the compression measure $C_{l}$ in the logarithmic scale; In the right figures, the $y$-axis denotes the discrimination measure $D_{l}$. More experimental details can be found in Section 5.2.2.
@@
Although our result is rooted in DLNs, it provides valuable insights into the feature evolution in nonlinear networks. Specifically, the linear separability of features learned by initial layers in nonlinear networks (see Figure 1) allows later layers to be effectively replaced by linear counterparts for compressing within-class features and discriminating between-class features. We empirically support this claim in Figure 6, where we observe that both the linear and subsequent nonlinear layers exhibit consistent trends of feature compression and discrimination with respect to depth. Therefore, studying DLNs helps us understand the role of the nonlinear layers after the initial layers in nonlinear networks for learning features. This understanding also sheds light on the pattern where within-class features compress at a geometric rate and between-class features discriminate at a linear rate in nonlinear networks, as illustrated in Figure 5 (bottom figures). A natural direction is to extend our current analysis framework to nonlinear networks, especially homogeneous neural networks (lyu2019gradient).
(a) A 8-layer MLP network with ReLU
(b) A 8-layer hybrid network
Figure 6: Within-class compression and between-class discrimination of features of two 8-layer networks. We use the same setup as in Figure 1 and plot the metrics of within-class compression and between-class discrimination (see Definition 1) of the two networks across layers after training, respectively. Additional experimental details are deferred to Section 5.1.1.